OpenStudy (anonymous):

Can somebody please explain to me how to solve for the midpoint & length of line segments? Thank you.

OpenStudy (anonymous):

Here is the question

OpenStudy (anonymous):

Any example is fine

OpenStudy (anonymous):

Midpoint = (x1+x2)/2, (y1+y2)/2

OpenStudy (anonymous):

I know the formula, but I'm unsure of how to apply it

OpenStudy (anonymous):

Length of line segments \[\sqrt{(y2-y1)^{2}+(x2-x1)^{2}}\]

OpenStudy (anonymous):

\[\frac{ -2 }{ 7 }=\frac{ -5 + x2 }{ -1 + y2 }\] Can you solve from there?

OpenStudy (anonymous):

Ah they gave you an end point and the mid point. That is so aggravating. You have to find the other end point using the formula y=m*x+b

OpenStudy (anonymous):

Okay I see, I think I've got it now! Much thanks to both of you.